A simple pendulum has a period of 512s. When the length is shorten by 2m, it’s period is 128s. Calculate the original length of the pendulum


1
Expert's answer
2021-01-15T05:13:19-0500
"T=2\pi \sqrt{\frac{l}{g}}\\
\frac{512}{128}= \sqrt{\frac{l}{l-2}}\\l=2.13\ m"